OpenStudy (anonymous):

help please What is the solution to r over three is equal to one over four? one over twelve twelve over one three over four four over three

OpenStudy (anonymous):

jeez whatever happened to arabic numerals?

OpenStudy (anonymous):

\[\frac{r}{3}=\frac{1}{4}\] maybe ?

OpenStudy (anonymous):

That's part of the question

OpenStudy (anonymous):

@lonnie455rich u like math? Well can u please help me out with this

OpenStudy (anonymous):

if \[\frac{r}{3}=\frac{1}{4}\] then multiply by \(3\) to get \[r=\frac{3}{4}\]

OpenStudy (anonymous):

who kicked you out?

OpenStudy (anonymous):

try now